OpenStudy (anonymous):

B^2+ab factored I had this question on a test and I'm trying to figure it out but my answer makes sense so I'm not sure please help

OpenStudy (anonymous):

you want to factorize b^2+ab right ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

well we have b^2=b*b so b^2-ab=bb-ab=b(b-a)

OpenStudy (anonymous):

is that making a sense to you ?

OpenStudy (anonymous):

Sort of

OpenStudy (anonymous):

what is step you didn't understand ?

OpenStudy (anonymous):

I'm not sure its just when try multiplying it back out it doesnt work

OpenStudy (anonymous):

let's try b(b-a)=b*b-b*a=b^2-ba

OpenStudy (anonymous):

Its b^2+ab not b^2-ab

OpenStudy (anonymous):

do you have in the question b^2+ab or b^2-ab ?

OpenStudy (anonymous):

b^2+ab